OpenStudy (anonymous):

A system of equations is shown below 5x - 5y = 10 3x - 2y = 2 Part A. Create and equivelant system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this Part B. Show that the equivelant system has the same solution as the original system of equations.

OpenStudy (anonymous):

multiply the top eq by 2 multiply the bottom eq by 5

OpenStudy (anonymous):

I'm still confused that doesn't exactly work

OpenStudy (jdoe0001):

hmm so.. have you covered "system of equations" yet?

OpenStudy (jdoe0001):

or... do you know how to solve a system of equations? have you covered either, substitution or elimination yet?

OpenStudy (anonymous):

Kind of but it still just really confuses me

OpenStudy (anonymous):

Can you explain to me how to solve the problem?

OpenStudy (jdoe0001):

"equivelant system of equations by replacing one equation with the sum of that equation and a multiple of the other" so \(\large { \begin{array}{rrrrrr} 5x - 5y = 10&\implies &5x - 5y = 10\\ 3x - 2y = 2 &{\color{brown}{ \times \square }}&\implies \square \cdot 3x+\square \cdot -2y=\square \cdot 10 \\\hline\\ \end{array} }\) so.. notice you'd grab either, in this case, we use the 1st equation as is and then you grab the 2nd one and multiply times "something", you pick what value then sum them up vertically and that'd give you a new equation, that you can use instead of either of the two original ones you had

OpenStudy (jdoe0001):

say.. pick a value.... any value that we could use to multiply the 2nd one

OpenStudy (anonymous):

I'm still not quite understanding sorry

OpenStudy (anonymous):

2

OpenStudy (jdoe0001):

ok.. so... let's start then by picking a value.. .ok so we'll use 2

OpenStudy (anonymous):

Ok

OpenStudy (jdoe0001):

\(\large {\begin{array}{crrcrrr} 5x - 5y = 10&&\implies &5x - 5y = 10\\ 3x - 2y = 2 &{\color{brown}{ \times 2 }}&\implies &6x-4y=4 \\\hline\\ &&&\textit{sum them up} \end{array} }\) notice.... you'd multiply, in this case the 2nd equation by 2 then, sum up, the result from the multiplication, and the other equation what would the sum give you?

OpenStudy (anonymous):

But how do I find the values for the variables?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

you'd sum them up vertically

OpenStudy (anonymous):

What does that mean

OpenStudy (jdoe0001):

hmmm well... you said you have solved system of equations already so.... is the same as what you used in doing the "elimination" method

OpenStudy (anonymous):

I have but I am not seeing how it relates

OpenStudy (anonymous):

I'm really sorry this is just really confusing me

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well... do you see the multiplication by 2 at least?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

ok.. how about the sum? not obvious enough? \(\large {\begin{array}{crrcrrr} 5x - 5y = 10&&\implies &{\color{purple}{5x - 5y = 10 }}\\ 3x - 2y = 2 &{\color{brown}{ \times 2 }}&\implies &{\color{purple}{ 6x-4y=4 }} \\\hline\\ &&&\textit{sum them up} \end{array} }\)

OpenStudy (anonymous):

I understand that

OpenStudy (jdoe0001):

ok... so... what would that sum give you though?

OpenStudy (anonymous):

Are you asking me to multiply the top equation by 2 as well @jdoe0001

OpenStudy (jdoe0001):

hmmm nope, to "sum them up"

OpenStudy (anonymous):

How do I do that @jdoe0001

OpenStudy (jdoe0001):

hmmm you may want to check your papers on the "elimination method" that should cover sum of the equations vertically

OpenStudy (anonymous):

II am in a quiz iso I can't otherwise I would

OpenStudy (anonymous):

11x + 1y= 6?? @jdoe0001

OpenStudy (anonymous):

I am not cheating this for something else not part of the quiz sorry if that was unclear

OpenStudy (anonymous):

@pooja195